OpenStudy (anonymous):

The value of \[6+\log_{\frac{ 3 }{ 2 }}(\frac{ 1 }{ 3\sqrt{2} }\sqrt{4-\frac{ 1 }{ 3\sqrt{2} }\sqrt{4-\frac{ 1 }{ 3\sqrt{2} }....}} \]

OpenStudy (anonymous):

Did you simplify the surd?

OpenStudy (anonymous):

Yup...) x= sqrt4-1/3sqrt2 sqrt4-1/3sqrt2...... squaring x^2 = 4 - 1/3sqrt2 *x

OpenStudy (anonymous):

You get, x = \[4 \sqrt2 /3\], Correct?

OpenStudy (anonymous):

^^ The other value of x is a negative one, so only this is viable.

OpenStudy (anonymous):

i got \[x=\frac{ 8 }{ 3\sqrt{2} }\]

OpenStudy (anonymous):

Let us go with ur x...)

OpenStudy (anonymous):

It's the same.

OpenStudy (anonymous):

Silly...yes....

OpenStudy (anonymous):

Nw wat to Do

OpenStudy (anonymous):

After that its simple. You've done all the hard work.

OpenStudy (anonymous):

\[\log_{\frac{ 3 }{ 2 }}(\frac{ 4 }{ 9 }) \]

OpenStudy (anonymous):

\[\log_{3/2}(2/3)^2 = - 2\log _{3/2}3/2 = -2\]

OpenStudy (anonymous):

Yup,,,,,).. thxxxx

OpenStudy (anonymous):

So you get 4. :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i want to study logarithms More.......Can....assist me..

OpenStudy (anonymous):

i mean...Do u knw any Gud videos on that Topic which i can get

OpenStudy (anonymous):

Khan acedemy. best option :)

OpenStudy (anonymous):

Thxxx.) again